OpenStudy (anonymous):

Solve: 10x + 15 = 13x Select the best answer from the choices provided. x = 2 x = 5 x = 8 x = 12

OpenStudy (uri):

Subtract 10 both sides @hanwingspan

OpenStudy (anonymous):

still Don't understand..........

OpenStudy (uri):

10x + 15 = 13x -10 -10 10x-10x = cancelled 15=3x 15/3=x x=15/3 simplify